In a school group of 45 students, 37 did the Biology, Physics and 20 Math tasks. a) What is the probability that a student did the Math or Biology assignment? b) What is the probability that he did not do any of the three tasks?
1
Expert's answer
2020-05-25T21:20:42-0400

A=biology, "P(A)=37\/45"

B=Physics, "P(B)=37\/45"

C=Math, "P(C)=20\/45"


a) "P(A+C)=P(A)+P(C)-P(AC)=
\frac{37}{45}+\frac{20}{45}-\frac{37}{45}\cdot\frac{20}{45}\approx 0.9012"


b) "P(\overline{ABC})=(1-P(A))(1-P(B))(1-P(C))=\frac{8}{45}\cdot \frac{8}{45}\cdot\frac{25}{45}\approx0.0176"